OpenStudy (anonymous):

Whats..? ÷ (16-8)= 4 whats the answer to where the ? is

OpenStudy (anonymous):

32

OpenStudy (anonymous):

32 (32 / 8 = 4)

OpenStudy (anonymous):

32

OpenStudy (anonymous):

8*4=32

OpenStudy (lalaly):

\[? =4 \times 8 =32\]

OpenStudy (anonymous):

let ? be x \[\frac{x}{16-8}=4\] \[\frac{x}{8}=4\] \[x=8*4\] \[x=32\]